A cricket player bowls a 156g cricket ball, releasing it from a height of exactly 1.1m. Ignoring the effect of air resistance, if the cricket player wants to release the ball such that its velocity is exactly parallel to the ground, at what velocity does she need to release it so that it will land 13.8m from the point of release?


Note: The height the ball has been released may have changed.


Give your answer to two decimal places in units of m/s. Do not include the units in your answer i.e. you would write 8.66 m/s as 8.66


1
Expert's answer
2021-03-24T19:41:18-0400

Let's first find the time that the ball takes to reach the ground:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot1.1\ m}{9.8\ \dfrac{m}{s^2}}}=0.47\ s."

Finally, we can find the necessary horizontal velocity:


"x=v_0t,""v_0=\dfrac{x}{t}=\dfrac{13.8\ m}{0.47\ s}=29.36\ \dfrac{m}{s}."